# Universal Digitized Counterdiabatic Driving

### v1

@@ Abstract
Introduction.— Adiabatic control is a fundamental means of tailoring quantum states with slow parameter changes [5, 25]. The adiabatic gauge potential plays an important role in the theory of adiabatic control [27]. It has various applications: it is a mathematical tool to prove the adiabatic theorem [25]; it is used to analyze near-adiabatic dynamics [33, 35, 18]; and it can be a probe to explore quantum phase transitions [43, 44] and quantum chaos [31, 28]. The adiabatic gauge potential is also central to the theory of shortcuts to adiabaticity, which enables to speed up adiabatic control [37, 17, 21]. Shortcuts to adiabaticity are promising candidates for improving quantum adiabatic algorithms, e.g., quantum annealing and quantum adiabatic computation [1], and thus have received much attention.
@@
Counterdiabatic driving is a method of shortcuts to adiabaticity using the adiabatic gauge potential [12, 11, 2]. Originally, the calculation of the adiabatic gauge potential required knowledge of energy spectra, but the theory of variational counterdiabatic driving reformulated it as a variational minimization problem without that knowledge [34]. This variational approach is based on the algebraic characterization of the adiabatic gauge potential, and thus it can also be reformulated as an algebraic problem [19, 41, 36, 4, 30]. More recently, universal counterdiabatic driving was proposed, based on universal polynomial fitting of the adiabatic gauge potential [29, 13]. These approaches theoretically enable to calculate the adiabatic gauge potential, but they require incredible computational cost because the number of relevant operators exponentially increases with system size in many-body systems.
@@
Truncation of many-body and/or nonlocal interactions in the adiabatic gauge potential is a typical approach to avoid expensive computational cost. Truncated adiabatic gauge potentials can be expressed in terms of nested commutators of a target Hamiltonian and its parameter derivative, and their coefficients can be numerically calculated with existing methods [34, 19, 41, 36, 4, 30, 29, 13]. Moreover, such terms can be effectively implemented using Floquet engineering [32, 10]. Besides Floquet engineering, digital quantum simulation is another approach to the realization [24, 23, 9, 20, 26, 38, 22, 39, 3]. These approaches are promising, but experimental realization is still limited up to a few-order nested commutators. A main obstacle in Floquet engineering is the requirement of high-frequency driving, and that in digital quantum simulation is the necessity of deep circuit depth.
@@
In this paper, we propose universal digitized counterdiabatic driving, a universal method of calculating the adiabatic gauge potential and realizing counterdiabatic driving in a digital way. We emphasize that it is not simple digitization of universal counterdiabatic driving, but it is a pivotal universal method of digitized counterdiabatic driving. Our method does not introduce any many-body and/or nonlocal interactions, which are not included in a target Hamiltonian. Nevertheless, we can digitally generate infinite nested commutators, which constitute the adiabatic gauge potential. Moreover, we can obtain explicit expression of rotation angles for digital implementation. In addition to these clear features, our analysis shows that the present method leads to the resolution of the previous problems mentioned above, making it promising.
@@
Theoretical background.— We briefly summarize the theory of counterdiabatic driving, from the definition of the adiabatic gauge potential to the idea of universal counterdiabatic driving. For a given parametrized Hamiltonian of a $D$-dimensional quantum system, $\hat{H}(\lambda)=\sum_{n=0}^{D-1}E_{n}(\lambda)|n(\lambda)\rangle\langle n(\lambda)|$, the adiabatic gauge potential is defined as
@@
The adiabatic gauge potential generates small displacement of an energy eigenstate in parameter space as $|n(\lambda+\delta\lambda)\rangle=\hat{U}_{\mathrm{ad}}(\lambda)|n(\lambda)\rangle$, where $\hat{U}_{\mathrm{ad}}(\lambda)$ is a unitary operator
@@
with infinitesimal parameter displacement $\delta\lambda$ [27]. Equation (2) is regarded as the quench limit of counterdiabatic driving. Our target is to construct the adiabatic gauge potential (1) for realizing the unitary operator (2) in a practical way.
It is possible to approximate the adiabatic gauge potential (1) with an Hermitian operator
@@
where $\mathcal{L}$ is a superoperator $\mathcal{L}\bullet=[\hat{H}(\lambda),\bullet]$. The coefficients $\{\alpha_{l}(\lambda)\}$ can be determined by using the algebraic characterization of the adiabatic gauge potential [34, 19, 41, 36, 4, 30] or polynomial fitting [29, 13]. It can actually be the exact adiabatic gauge potential (1) when the integer $d$ is large and the nested commutators $\{\mathcal{L}^{2l-1}\partial_{\lambda}\hat{H}(\lambda)\}$ constitute closed Lie algebra. Owing to the difficulties in computational cost and experimental implementation, small $d$, usually $d=\mathcal{O}(1)$, is adopted in practice.
The difference between the Hermitian operator (3) and the adiabatic gauge potential (1) can be evaluated as
@@
Here, $\mathrm{Si}(x)=\int_{0}^{x}dy(1/y)\sin(y)$ is the sine integral and $\Omega$ is a certain cutoff. As discussed later, it is natural to set $\Omega=\Delta_{\mathrm{max}}$, where $\Delta_{\mathrm{max}}$ is the maximum energy gap, while it is not necessarily to set in that way. Clearly, no many-body and/or nonlocal interactions, which are not included in the target Hamiltonian $\hat{H}(\lambda)$, are introduced.
We explain the key ideas and advantages of our method (details of the derivation are available at END MATTER). The composite unitary operator (5) can be approximated as a unitary operator with a single generator, i.e., $\hat{U}(\lambda)=e^{-i\delta\lambda\hat{V}(\lambda)}+\mathcal{O}(\{\phi_{k}^{2}\})$ with an Hermitian operator
@@
for $\theta_{k}=-\theta_{-k}$, $\phi_{k}=-\phi_{-k}$, and $|\phi_{k}|\ll 1$. The operator (7) has similar structure to the operator (3), and thus it has the potential to approximate the adiabatic gauge potential (1). Notably, the operator (7) consists of infinite nested commutators unlike the operator (3).
The difference between the operator (7) and the adiabatic gauge potential (1) is given by
@@
Remarkably, owing to the infinite nested commutators in Eq. (7), we can adopt the Fourier series expansion instead of polynomial fitting unlike universal counterdiabatic driving, and it enables to obtain the explicit expression of the rotation angles. Indeed, the Fourier expansion of $(-1/\omega)$ gives the rotation angles (6). The spectral function $\Phi(\omega)$ is nonzero only in the ranges $\Delta_{\mathrm{max}},-\Delta_{\mathrm{min}}]$ and $\Delta_{\mathrm{min}},\Delta_{\mathrm{max}}]$, where $\Delta_{\mathrm{min}}$ is the minimum energy gap. This is the reason why $\Omega=\Delta_{\mathrm{max}}$ is natural, while we will numerically show that it is not necessarily to set in that way. Note that, when the target state is the ground state, $\Delta_{\mathrm{min}}$ can be regarded as the minimum energy gap for the ground state, not the minimum energy gap between all the eigenstates. In summary, the validity of universal digitized counterdiabatic driving, using the composite unitary operator (5) with the rotation angles (6), is theoretically supported.
@@
First, we consider a two-level system $\hat{H}(\lambda)=h^{X}(\lambda)\hat{X}+h^{Z}(\lambda)\hat{Z}$ with parametrized fields $\{h^{X}(\lambda),h^{Z}(\lambda)\}$, where $\{\hat{X},\hat{Y},\hat{Z}\}$ are the Pauli matrices. The energy gap of this system is given by $\Delta=2\sqrt{h^{X}(\lambda)^{2}+h^{Z}(\lambda)^{2}}$ and the spectral function $\Phi(\omega)$ is nonzero only for $\omega=\pm\Delta$. Thus, the difference (8) becomes exactly zero for rotation angles $\theta_{1}=\pi/2\Delta$ and $\phi_{1}=-\delta\lambda/\Delta$ with $K=1$. For these rotation angles, the operator (7) is actually identical with the exact adiabatic gauge potential of the two-level Hamiltonian [12, 11, 2] and the unitary operator (5) approximates counterdiabatic driving (2) as $\hat{U}(\lambda)=\hat{U}_{\mathrm{ad}}(\lambda)+\mathcal{O}(\delta\lambda^{2})$ (see END MATTER for more details). Therefore, the basics of our method are consistent with the exact theory.
@@
Next, we consider a many-body system $\hat{H}(\lambda)=[J(\lambda)/2N]\sum_{i,j=1}^{N}\hat{Z}_{i}\hat{Z}_{j}+h^{X}(\lambda)\sum_{i=1}^{N}\hat{X}_{i}$ with parameters $\{J(\lambda),h^{X}(\lambda)\}$, where $\{\hat{X}_{i},\hat{Y}_{i},\hat{Z}_{i}\}_{i=1}^{N}$ are the Pauli matrices of $N$ spins. The ground state of this system shows the second-order quantum phase transition at the critical point $|h^{X}(\lambda)/J(\lambda)|=1$ for $J(\lambda)<0$ [6, 7, 8, 42]. The system Hamiltonian commutes with the total-spin operator $\sum_{W=X,Y,Z}(\sum_{i=1}^{N}\hat{W_{i}})^{2}$ and the parity operator $\prod_{i=1}^{N}\hat{X}_{i}$, and thus dynamics is confined in $(N+1)$ dimension when the initial state is symmetric for these two operators. The ground state satisfies this condition.
@@
We numerically study parameter displacement of the ground state $|0(\lambda)\rangle$ at the critical point with $J(\lambda)=J_{0}=-1$, $h^{X}(\lambda)=h_{0}^{X}\lambda=1$ ($h_{0}^{X}=1$ and $\lambda=1$), $\delta\lambda=10^{-3}$, and $N=10$. For our method, we set $\Omega=\Delta_{\mathrm{max}}=20.278\cdots$ and adopt the rotation angles (6). The infidelity to the ground state is plotted with respect to the integer $K$ in Fig. 1.
@@
Figure 1: The infidelity to the ground state of the many-body system with respect to the integer $K$. The parameters are $J(\lambda)=J_{0}=-1$, $h^{X}(\lambda)=h_{0}^{X}\lambda=1$ ($h_{0}^{X}=1$ and $\lambda=1$), $\delta\lambda=10^{-3}$, $N=10$, and $\Omega=\Delta_{\mathrm{max}}=20.278\cdots$. The black solid line represents the infidelity under the parameter quench and the black dashed lines represent $kK_{p}^{\Omega=\Delta_{\mathrm{max}}}/2$ with $k=1,2,\dots$.
@@
We find the success in achieving very low infidelity in a periodic way with a period $K_{p}^{\Omega=\Delta_{\mathrm{max}}}=\Delta_{\mathrm{max}}/\Delta_{\mathrm{min}}$ (the black dashed lines represent $kK_{p}^{\Omega=\Delta_{\mathrm{max}}}/2$ with $k=1,2,\dots$), while the infidelity returns to the original value periodically (the black solid line represents the infidelity under a parameter quench, i.e., $1-|\langle 0(\lambda+\delta\lambda)|0(\lambda)\rangle|^{2}$).
@@
We then analyze this periodic behavior. We expect that the infidelity becomes finite mainly because of the difference of the operator (7) from the exact adiabatic gauge potential (1), and thus we focus on the kernel of the difference (8), i.e., $(1/\omega)+\sum_{k=1}^{K}(\phi_{k}/\delta\lambda)\sin(\theta_{k}\omega)]^{2}$. It is known that the sine integral converges to $\pi/2$, i.e., $\lim_{x\to\infty}\mathrm{Si}(x)=\pi/2$, and accordingly $\phi_{k}$ also converges to $(-\pi\delta\lambda/\Omega)$ for large $k$. This means that the difference (8) does not converge to zero because the summand converges to $(-\pi/\Omega)\sin(k\pi\omega/\Omega)$, associated with the Gibbs phenomenon [40, 14, 15] of the Fourier series due to the singularity of $(1/\omega)$ at $\omega=0$. By setting $k\to k^{\prime}K_{p}^{\Omega=\Delta_{\mathrm{max}}}/2$ ($k^{\prime}=1,2,\dots$) and $\omega=\Delta_{\mathrm{min}}$, the summand becomes $(-\pi/\Delta_{\mathrm{max}})\sin(k^{\prime}\pi/2)$. That is, when oscillation of the summand is canceled out, which is achieved for odd $k^{\prime}$, we can obtain very low infidelity. We also plot the kernel for several $K$ in Fig. 2.
@@
Figure 2: Difference from the exact adiabatic gauge potential with respect to the energy differences. Each curve represents (red solid curve) $K=4$, (green dashed curve) $K=8$, (blue dotted curve) $K=13$, and (yellow dashed-dotted curve) $K=17$. The parameters are the same as Fig. 1. The gray shaded areas are out of $\Delta_{\mathrm{min}},\Delta_{\mathrm{max}}]$. The inset is an enlarged view around the minimum energy gap $\Delta_{\mathrm{min}}$.
@@
The gray shaded areas are out of $\Delta_{\mathrm{min}},\Delta_{\mathrm{max}}]$, and thus it is important to suppress the kernel in the white area, particularly around the minimum energy gap. We actually find the cancellation of the summand for $K=4,13$ (corresponding to $k^{\prime}=1,3$) and accumulation of it for $K=8,17$ (corresponding to $k^{\prime}=2,4$).
@@
Finally, we show that the cutoff $\Omega$ is not necessarily the maximum energy gap $\Delta_{\mathrm{max}}$ because the highest-energy eigenstate is irrelevant to the ground state. To exemplify this, we plot the infidelity to the ground state similarly to Fig. 1 with the cutoff $\Omega=15$ instead of the maximum energy gap $\Delta_{\mathrm{max}}=20.278\cdots$ in Fig. 3.
@@
Figure 3: The infidelity to the ground state of the many-body system with respect to the integer $K$. The parameters are the same as Fig. 1 except for the cutoff $\Omega=15$ instead of the maximum energy gap $\Delta_{\mathrm{max}}=20.278\cdots$.
@@
Discussion.— Now we compare our method with previously-proposed methods. The most relevant digital method was proposed in Ref. [38] and benchmarked in Ref. [22], which also used alternating unitary operators like Eq. (5). However, this method was derived by introducing several approximations to Eq. (2), and thus control fidelity was very low and results were very noisy because of several approximation parameters. In contrast, our method has only two free parameters: the integer $K$ and the cutoff $\Omega$. Remarkably, our method has achieved very low infidelity, and the behavior of our method is periodic against the integer $K$ and robust against the cutoff $\Omega$. That is, our method is stable.
More recently, another digital method was proposed in Ref. [3]. In this method, a unitary operator of the truncated adiabatic gauge potential (3) is systematically decomposed into alternating unitary operators by using product formulae. Main disadvantages of this method are the requirement of the exponentially large number of unitary operators for composing nested commutators and the necessity of determining rotation angles with numerical calculation. In contrast, our method does not necessarily impose an exponential realization cost, and gives the explicit expression of the rotation angles. Our method also includes convincingly-controlled infinite nested commutators unlike their method.
@@
Our method has addressed the problems of the previous methods mentioned in the introduction: explicit expression of the unitary operator and the rotation angles is given, i.e., there is no expensive computational cost; convincingly-controlled infinite nested commutators are included; and the requirement of high-frequency driving or deep circuit depth can be mitigated. A remaining limitation of our method is large $K$ required for the (exponentially) small minimum energy gap $\Delta_{\mathrm{min}}$ as well as other methods.
Conclusion and outlook.— In this paper, we have proposed universal digitized counterdiabatic driving, a universal method of constructing the adiabatic gauge potential and realizing counterdiabatic driving in a digital way. Our method does not require additional many-body and/or nonlocal interactions; includes convincingly-controlled infinite nested commutators; and gives the explicit expression of the rotation angles. We have shown that our method can achieve very low infidelity even in a many-body system.
An important direction extending the present work is the development of methods suppressing deviations around the minimum energy gap more efficiently. One of the approaches would be the regularization of the adiabatic gauge potential removing the divergence [10], while it requires knowledge of the minimum energy gap to approximate the adiabatic gauge potential well. Standard approaches in the literature of the Fourier analysis, e.g., the use of a window function for suppressing the Gibbs phenomenon [16], are also applicable. Thus, we could extend our method furthermore. In our demonstration, we just focus on the parameter displacement of the energy eigenstates, but the adiabatic gauge potential has various applications [25, 33, 35, 18, 43, 44, 31, 28] as mentioned in the introduction. Developing our method for those applications is the important future work to be addressed.

### v2

@@ Abstract
Introduction.— Adiabatic control is a fundamental means of tailoring quantum states with slow parameter changes [5, 26]. The adiabatic gauge potential (AGP) plays an important role in the theory of adiabatic control [28]. It has various applications: it is a mathematical tool to prove the adiabatic theorem [26]; it is used to analyze near-adiabatic dynamics [34, 36, 18]; and it can be a probe to explore quantum phase transitions [44, 45] and quantum chaos [32, 29]. AGP is also central to the theory of shortcuts to adiabaticity, which enables to speed up adiabatic control [38, 17, 21]. Shortcuts to adiabaticity are promising candidates for improving quantum adiabatic algorithms, e.g., quantum annealing and quantum adiabatic computation [1], and thus have received much attention.
@@
Counterdiabatic driving is a method of shortcuts to adiabaticity using AGP [12, 11, 2]. Originally, the calculation of AGP required knowledge of energy spectra, but the theory of variational counterdiabatic driving reformulated it as a variational minimization problem without that knowledge [35]. This variational approach is based on the algebraic characterization of AGP, and thus it can also be reformulated as an algebraic problem [19, 42, 37, 4, 31]. More recently, universal counterdiabatic driving was proposed, based on universal polynomial fitting of AGP [30, 13]. These approaches theoretically enable to calculate AGP, but they require incredible computational cost because the number of relevant operators exponentially increases with system size in many-body systems.
@@
Truncation of many-body and/or nonlocal interactions in AGP is a typical approach to avoid expensive computational cost. Truncated AGPs can be expressed in terms of nested commutators of a target Hamiltonian and its parameter derivative [10], and their coefficients can be numerically calculated with existing methods [35, 19, 42, 37, 4, 31, 30, 13]. Moreover, such terms can be effectively implemented using Floquet engineering [33, 10]. Besides Floquet engineering, digital quantum simulation is another approach to the realization [25, 24, 9, 20, 27, 39, 22, 40, 3]. These approaches are promising, but experimental realization is still limited up to a few-order nested commutators. A main obstacle in Floquet engineering is the requirement of high-frequency driving, and that in digital quantum simulation is the necessity of deep circuit depth.
@@
In this paper, we propose universal digitized counterdiabatic driving, a universal method of calculating AGP and realizing counterdiabatic driving in a digital way. We emphasize that it is not simple digitization of universal counterdiabatic driving, but it is a pivotal universal method of digitized counterdiabatic driving. Our method does not introduce any many-body and/or nonlocal interactions, which are not included in a target Hamiltonian. Nevertheless, we can digitally generate infinite nested commutators, which constitute AGP. Moreover, we can obtain explicit expression of rotation angles for digital implementation. In addition to these clear features, our analysis shows that the present method leads to the resolution of the previous problems mentioned above, making it promising.
@@
Theoretical background.— We briefly summarize the theory of counterdiabatic driving, from the definition of AGP to the idea of universal counterdiabatic driving. For a given parametrized Hamiltonian of a $D$-dimensional quantum system, $\hat{H}(\lambda)=\sum_{n=0}^{D-1}E_{n}(\lambda)|n(\lambda)\rangle\langle n(\lambda)|$, AGP is defined as
@@
AGP generates small displacement of an energy eigenstate in parameter space as $|n(\lambda+\delta\lambda)\rangle=\hat{U}_{\mathrm{ad}}(\lambda)|n(\lambda)\rangle$, where $\hat{U}_{\mathrm{ad}}(\lambda)$ is a unitary operator
@@
with infinitesimal parameter displacement $\delta\lambda$ [28]. Equation (2) is regarded as the quench limit of counterdiabatic driving. Our target is to construct AGP (1) for realizing the unitary operator (2) in a practical way.
It is possible to approximate AGP (1) with an Hermitian operator
@@
where $\mathcal{L}$ is a superoperator $\mathcal{L}\bullet=[\hat{H}(\lambda),\bullet]$ [10]. The coefficients $\{\alpha_{l}(\lambda)\}$ can be determined by using the algebraic characterization of AGP [35, 19, 42, 37, 4, 31] or polynomial fitting [30, 13]. It can actually be the exact AGP (1) when the integer $d$ is large and the nested commutators $\{\mathcal{L}^{2l-1}\partial_{\lambda}\hat{H}(\lambda)\}$ constitute closed Lie algebra. Owing to the difficulties in computational cost and experimental implementation, small $d$, usually $d=\mathcal{O}(1)$, is adopted in practice.
The difference between the Hermitian operator (3) and AGP (1) can be evaluated as
@@
Here, $\mathrm{Si}(x)=\int_{0}^{x}dy(1/y)\sin(y)$ is the sine integral and $\Omega$ is a certain cutoff. As discussed later, it is natural to set $\Omega=\Delta_{\mathrm{max}}$, where $\Delta_{\mathrm{max}}$ is the maximum energy gap, while it is not necessarily to set in that way.
We explain the key ideas, advantages, and possible extension of our method (details of the derivation are available at END MATTER). Clearly, the composite unitary operator (5) does not have any many-body and/or nonlocal interactions, which are not included in the target Hamiltonian $\hat{H}(\lambda)$. The composite unitary operator (5) can be approximated as a unitary operator with a single generator, i.e., $\hat{U}(\lambda)=e^{-i\delta\lambda\hat{V}(\lambda)}+\mathcal{O}(\{\phi_{k}^{2}\})$ with an Hermitian operator
@@
for $\theta_{k}=-\theta_{-k}$, $\phi_{k}=-\phi_{-k}$, and $|\phi_{k}|\ll 1$. The operator (7) has similar structure to the operator (3), and thus it has the potential to approximate AGP (1). Notably, the operator (7) consists of infinite nested commutators unlike the operator (3).
The difference between the operator (7) and AGP (1) is given by
@@
Remarkably, owing to the infinite nested commutators in Eq. (7), we can adopt the Fourier series expansion instead of polynomial fitting unlike universal counterdiabatic driving, and it enables to obtain the explicit expression of the rotation angles. Indeed, the Fourier expansion of $(-1/\omega)$ gives the rotation angles (6). The spectral function $\Phi(\omega)$ is nonzero only in the ranges $\Delta_{\mathrm{max}},-\Delta_{\mathrm{min}}]$ and $\Delta_{\mathrm{min}},\Delta_{\mathrm{max}}]$, where $\Delta_{\mathrm{min}}$ is the minimum energy gap. This is the reason why $\Omega=\Delta_{\mathrm{max}}$ is natural, while we will numerically show that it is not necessarily to set in that way. In summary, the validity of universal digitized counterdiabatic driving, using the composite unitary operator (5) with the rotation angles (6), is theoretically supported.
A possible extension of our proposal is the introduction of the regularization in AGP [10], which results in $(1/\omega)\to\omega/(\omega^{2}+\eta^{2})$ ($\eta^{2}\ll\Delta_{\mathrm{min}}^{2}$) in Eq. (8). It resolves a certain problem associated with the divergence of $(1/\omega)$ at $\omega=0$, which we will find by numerical simulation, while $\phi_{k}$ is modulated as $\phi_{k}=-\mathrm{sgn}(k)\frac{2\delta\lambda}{\Omega}\int_{0}^{\Omega}[\omega/(\omega^{2}+\eta^{2})]\sin(k\pi\omega/\Omega)d\omega$. Note that, when the target state is the ground state, $\Delta_{\mathrm{min}}$ can be regarded as the minimum energy gap for the ground state, not the minimum energy gap between all the eigenstates.
@@
First, we consider a two-level system $\hat{H}(\lambda)=h^{X}(\lambda)\hat{X}+h^{Z}(\lambda)\hat{Z}$ with parametrized fields $\{h^{X}(\lambda),h^{Z}(\lambda)\}$, where $\{\hat{X},\hat{Y},\hat{Z}\}$ are the Pauli matrices. The energy gap of this system is given by $\Delta=2\sqrt{h^{X}(\lambda)^{2}+h^{Z}(\lambda)^{2}}$ and the spectral function $\Phi(\omega)$ is nonzero only for $\omega=\pm\Delta$. Thus, the difference (8) becomes exactly zero for rotation angles $\theta_{1}=\pi/2\Delta$ and $\phi_{1}=-\delta\lambda/\Delta$ with $K=1$. For these rotation angles, the operator (7) is actually identical with the exact AGP of the two-level Hamiltonian [12, 11, 2] and the unitary operator (5) approximates counterdiabatic driving (2) as $\hat{U}(\lambda)=\hat{U}_{\mathrm{ad}}(\lambda)+\mathcal{O}(\delta\lambda^{2})$ (see END MATTER for more details). Therefore, the basics of our method are consistent with the exact theory.
@@
Next, we consider a many-body system $\hat{H}(\lambda)=[J(\lambda)/2N]\sum_{i,j=1}^{N}\hat{Z}_{i}\hat{Z}_{j}+h^{X}(\lambda)\sum_{i=1}^{N}\hat{X}_{i}$ with parameters $\{J(\lambda),h^{X}(\lambda)\}$, where $\{\hat{X}_{i},\hat{Y}_{i},\hat{Z}_{i}\}_{i=1}^{N}$ are the Pauli matrices of $N$ spins. The ground state of this system shows the second-order quantum phase transition at the critical point $|h^{X}(\lambda)/J(\lambda)|=1$ for $J(\lambda)<0$ [6, 7, 8, 43]. The system Hamiltonian commutes with the total-spin operator $\sum_{W=X,Y,Z}(\sum_{i=1}^{N}\hat{W_{i}})^{2}$ and the parity operator $\prod_{i=1}^{N}\hat{X}_{i}$, and thus dynamics is confined in $\lceil(N+1)/2\rceil$ dimension when the initial state is symmetric for these two operators. The ground state satisfies this condition.
@@
We numerically study parameter displacement of the ground state $|0(\lambda)\rangle$ at the critical point with $J(\lambda)=J_{0}=-1$, $h^{X}(\lambda)=h_{0}^{X}\lambda=1$ ($h_{0}^{X}=1$ and $\lambda=1$), $\delta\lambda=10^{-3}$, and $N=10$. For our method, we set $\Omega=\Delta_{\mathrm{max}}=20.278\cdots$. The infidelity to the ground state is plotted with respect to the integer $K$ in Fig. 1 [the red circles are the results with the rotation angles (6) and the green squares are those with the regularization for $\eta=0.1\Delta_{\mathrm{min}}$].
@@
Figure 1: The infidelity to the ground state of the many-body system with respect to the integer $K$. The red circles are the results with the rotation angles (6) and the green squares are those with the regularization for $\eta=0.1\Delta_{\mathrm{min}}$. The parameters are $J(\lambda)=J_{0}=-1$, $h^{X}(\lambda)=h_{0}^{X}\lambda=1$ ($h_{0}^{X}=1$ and $\lambda=1$), $\delta\lambda=10^{-3}$, $N=10$, and $\Omega=\Delta_{\mathrm{max}}=20.278\cdots$. The black solid line represents the infidelity under the parameter quench and the black dashed lines represent $kK_{p}^{\Omega=\Delta_{\mathrm{max}}}/2$ with $k=1,2,\dots$.
@@
We find the success in achieving very low infidelity in a periodic way with a period $K_{p}^{\Omega=\Delta_{\mathrm{max}}}=\Delta_{\mathrm{max}}/\Delta_{\mathrm{min}}$ (the black dashed lines represent $kK_{p}^{\Omega=\Delta_{\mathrm{max}}}/2$ with $k=1,2,\dots$), while the infidelity returns periodically (the black solid line represents the infidelity under a parameter quench, i.e., $1-|\langle 0(\lambda+\delta\lambda)|0(\lambda)\rangle|^{2}$). Notably, the regularization can suppress the returns to the original value.
@@
We then analyze this periodic behavior with the rotation angles (6). We focus on the kernel of the difference (8), i.e., $(1/\omega)+\sum_{k=1}^{K}(\phi_{k}/\delta\lambda)\sin(\theta_{k}\omega)]^{2}$. It is known that the sine integral converges to $\pi/2$, i.e., $\lim_{x\to\infty}\mathrm{Si}(x)=\pi/2$, and accordingly $\phi_{k}$ also converges to $(-\pi\delta\lambda/\Omega)$ for large $k$. This means that the difference (8) does not converge to zero because the summand converges to $(-\pi/\Omega)\sin(k\pi\omega/\Omega)$, associated with the Gibbs phenomenon [41, 14, 15] of the Fourier series due to the singularity of $(1/\omega)$ at $\omega=0$. By setting $k\to k^{\prime}K_{p}^{\Omega=\Delta_{\mathrm{max}}}/2$ ($k^{\prime}=1,2,\dots$) and $\omega=\Delta_{\mathrm{min}}$, the summand becomes $(-\pi/\Delta_{\mathrm{max}})\sin(k^{\prime}\pi/2)$. That is, when oscillation of the summand is canceled out, which is achieved for odd $k^{\prime}$, we can obtain very low infidelity. We also plot the kernel for several $K$ in Fig. 2.
@@
Figure 2: The kernel that describes deviations from the exact AGP with respect to the energy differences. Each curve represents (red solid curve) $K=4$, (green dashed curve) $K=8$, (blue dotted curve) $K=13$, and (yellow dashed-dotted curve) $K=17$. The parameters are the same as Fig. 1. The gray shaded areas are out of $\Delta_{\mathrm{min}},\Delta_{\mathrm{max}}]$. The inset is an enlarged view around the minimum energy gap $\Delta_{\mathrm{min}}$.
@@
The gray shaded areas are out of $\Delta_{\mathrm{min}},\Delta_{\mathrm{max}}]$, and thus it is important to suppress the kernel in the white area, particularly around the minimum energy gap. We actually find the cancellation of the summand for $K=4,13$ (corresponding to $k^{\prime}=1,3$) and accumulation of it for $K=8,17$ (corresponding to $k^{\prime}=2,4$). Note that the regularization removes the singularity of $(1/\omega)$ at $\omega=0$, and thus it resolves the Gibbs phenomenon.
@@
Finally, we show that the cutoff $\Omega$ is not necessarily the maximum energy gap $\Delta_{\mathrm{max}}$ because the highest-energy eigenstate is irrelevant to the ground state. To exemplify this, we plot the infidelity to the ground state similarly to Fig. 1 except for the cutoff $\Omega=15$ instead of the maximum energy gap $\Delta_{\mathrm{max}}=20.278\cdots$ in Fig. 3.
@@
Figure 3: The infidelity to the ground state of the many-body system with respect to the integer $K$. The symbols and parameters are the same as Fig. 1 except for the cutoff $\Omega=15$ instead of the maximum energy gap $\Delta_{\mathrm{max}}=20.278\cdots$.
@@
Discussion.— Now we compare our method with previously-proposed methods. The most relevant digital method was proposed in Ref. [39] and benchmarked in Ref. [22], which also used alternating unitary operators like Eq. (5). However, this method was derived by introducing several approximations to Eq. (2), and thus control fidelity was not very high and results were very noisy because of several approximation parameters. In contrast, our method has only two free parameters: the integer $K$ and the cutoff $\Omega$. Remarkably, our method has achieved very low infidelity, and the behavior of our method is periodic against the integer $K$ and robust against the cutoff $\Omega$. That is, our method is stable. Moreover, the complexity of our method is $\mathcal{O}(\Delta_{\mathrm{min}}^{-1})$ (see END MATTER for details), while that of their method is $\mathcal{O}(\Delta_{\mathrm{min}}^{-3})$.
More recently, another digital method was proposed in Ref. [3]. In this method, a unitary operator of the truncated AGP (3) is systematically decomposed into alternating unitary operators by using product formulae. A main disadvantage of this method is the requirement of the exponentially large number of unitary operators for composing nested commutators. In contrast, our method does not necessarily impose an exponential realization cost unless the minimum energy gap closes in an exponential way, while our method includes convincingly-controlled infinite nested commutators unlike their method. Moreover, our method gives the explicit expression of the rotation angles, while their method necessitates numerical calculation with the existing methods [35, 19, 42, 37, 4, 31, 30, 13].
@@
Our method has addressed the problems of the previous methods mentioned in the introduction: the unitary operator and the rotation angles are explicitly given without expensive computational cost; convincingly-controlled infinite nested commutators are generated, i.e., truncation is not necessary; and the requirement of high-frequency driving or deep circuit depth can be mitigated. A remaining limitation of our method is large $K$ required for the exponentially-small minimum energy gap $\Delta_{\mathrm{min}}$ as well as other methods.
Conclusion and outlook.— In this paper, we have proposed universal digitized counterdiabatic driving, a universal method of constructing AGP and realizing counterdiabatic driving in a digital way. Our method does not require additional many-body and/or nonlocal interactions; includes convincingly-controlled infinite nested commutators; and gives the explicit expression of the rotation angles. We have shown that our method can achieve very low infidelity.
An open problem we leave is the way of determining optimal $K$ without explicit knowledge of the minimum energy gap. Another important direction extending the present work is the development of methods suppressing deviations around the minimum energy gap more efficiently and more stably. For example, standard approaches in the literature of the Fourier analysis, e.g., the use of a window function for suppressing the Gibbs phenomenon [16], are applicable. We expect that further improvement of our method could be possible by incorporating such techniques. In our demonstration, we just focus on the parameter displacement of the energy eigenstates, but AGP has various applications [26, 34, 36, 18, 44, 45, 32, 29] as mentioned in the introduction. Developing our method for those applications is the important future work to be addressed.